One person has 2 child.At least one of them was a girl.Then the Probability that 2 of the are girl is_____

Grade:11

2 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

 

Dear Sairam,

 

Prob of a boy is 1/2 and girl is 1/2

according to ques.  atleast one child is girl and other is also girl so

Prob = 1 * 1/2 = 1/2

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay

Sudheesh Singanamalla
13 years ago

A Person can have a boy or a girl as a child

so therefore

sample space = { BB ,BG,GB,GG }

 

therefore probabilty that both are girls is 1/4

 

approve the answer if its right